OpenStudy (anonymous):

Solve /8f+3/=5 for f

OpenStudy (unklerhaukus):

do you mean \[|8f+3|=5\]

OpenStudy (anonymous):

yes. sorry

OpenStudy (unklerhaukus):

take the two cases \[+(8f+3)=5\qquad (1)\] \[-(8f+3)=5\qquad (2)\]

OpenStudy (anonymous):

okay. then do I just solve normally?

OpenStudy (unklerhaukus):

yes solve each case as you normally would

OpenStudy (anonymous):

ok. i did and then i got f= 1/4 and f=-1

OpenStudy (unklerhaukus):

Great work

OpenStudy (anonymous):

thanks for your help!